ACTIVITY: fnfidng the PD

  1. to carry how much charge between two points having potential difference equal to 220V, 1760J of work is done?
  2. the EMF of a cell falls from 3 volts t 2.8 volts when its terminals are joined to n electrical load of 4 Ohms. calculate the internal resistance of the cell.
1
Expert's answer
2021-05-03T18:37:16-0400

potential difference = work done / charge

thus charge =work done / potential difference

=1760/220=8C


"\frac{e-V}{1} × \frac{R}{V}=r"

"\frac{3-2.8}{1} × \frac{4}{2.8} =0.29ohms"